OpenStudy (anonymous):

The answer to 7 + 4x = 13 + x

OpenStudy (anonymous):

Put the x's (4x, x) on one side anD the numbers on the other (7,13) of the equal sign. If you throw over a positive number, it becomes negative and vice versa. If you throw over a number that is multiplied, it divides. So, 1) 4x=13+x-7 2) 4x-1x=6 (1x because one x = 1x) 3) 3x=6 4) x=6/3 5) x=2